OpenStudy (anonymous):

4x^2+ 32x + 64=25 Solve the equation by using the Square Root Property. could someone help me

OpenStudy (anonymous):

I know how to solve, but what is a Square Root Property?

OpenStudy (anonymous):

is one method that is used to find the solutions to a quadratic (second degree) equation

OpenStudy (anonymous):

I see, it's actually called completing the square, but, anyway... factor out of 4, and divide both sides by 4. tell me what you get.

OpenStudy (anonymous):

Actually that might not be necessary.

OpenStudy (anonymous):

\[(2x+16)^2=25\]

OpenStudy (anonymous):

\[2x+16=±\sqrt{25}\]\[2x+16=±5\]\[2x=-16±5\]\[x=\frac{-16±5}{2}\]\[So,\]\[x=-\frac{21}{2}\]\[and\]\[x=-\frac{11}{2}\]